Suppose that we will take a random sample of size n from a population having mean u and standard deviation o. For each of



the following situations, find the mean, variance, and standard deviation of the sampling distribution of the sample mean X:




a. μ = 10, o=2, n = 25



b. μ=500, o=.5, n = 100



C. μ=3, o =., n=4



d. μ= 100, o=1, n = 1,600




What will I do here if the mean and stdv is already given? TT sorry I am confused

1
Expert's answer
2022-05-02T15:15:18-0400

a.

"\mu_{\bar{X}}=\mu=10"

"Var(\bar{X})=\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}=\dfrac{2^2}{25}=0.16"

"\sigma_{\bar{X}}=\dfrac{\sigma}{\sqrt{n}}=\dfrac{2}{\sqrt{25}}=0.4"

b.

"\mu_{\bar{X}}=\mu=500"

"Var(\bar{X})=\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}=\dfrac{0.5^2}{100}=0.0025"

"\sigma_{\bar{X}}=\dfrac{\sigma}{\sqrt{n}}=\dfrac{0.5}{\sqrt{100}}=0.05"

c.

"\mu_{\bar{X}}=\mu=3"

"Var(\bar{X})=\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}=\dfrac{0.1^2}{4}=0.0025"

"\sigma_{\bar{X}}=\dfrac{\sigma}{\sqrt{n}}=\dfrac{0.1}{\sqrt{4}}=0.05"

d.

"\mu_{\bar{X}}=\mu=100"

"Var(\bar{X})=\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}=\dfrac{1^2}{1600}=0.000625"

"\sigma_{\bar{X}}=\dfrac{\sigma}{\sqrt{n}}=\dfrac{1}{\sqrt{1600}}=0.025"